PURE ROTATIONAL SPECTROSCOPY OF PANHs I: 1,10-PHENANTHROLINE. IMPLICATIONS OF PANHs IN ASTROPHYSICAL ENVIRONMENTS AND OBSERVATIONAL SPECTRA

\underline{BRETT A. MCGUIRE}, IAN A. FINNERAN, P. BRANDON CARROLL, Department of Chemistry, California Institute of Technology, Pasadena, CA 91125; and GEOFFREY A. BLAKE, Divisions of Geological & Planetary Sciences and Chemistry & Chemical Engineering, California Institute of Technology, Pasadena, CA 91125.

Polycyclic aromatic hydrocarbons (PAHs) have been proposed as possible carriers of the diffuse interstellar bands, the unidentified infrared features, and as likely precursors for the recently-observed C_{60} and C_{70} fullerenes. While PAHs have been well studied in the laboratory, and ultraviolet through infrared spectral simulations of PAHs can reproduce astronomical spectra reasonably well, several discrepancies still exist. Nitrogen-substituted PAHs, PANHs, have been proposed as a possible explanation for one of the major differences: the peak position of the 6.2 μm feature. While identification of individual PAH and PANH species from infrared spectra alone is extremely difficult, identification in the mm and sub-mm regimes using heterodyne spectroscopy is far more feasible. The frequently low (or zero) dipole moment of PAHs makes pure-rotational laboratory measurements and astronomical observation difficult. PANHs, however, often have substantial dipole moments, making them ideal targets for laboratory and astronomical studies. We present here the results of a laboratory study of the PANH 1,10-phenanthroline using direct absorption mm/sub-mm spectroscopy. We discuss implications of these results for the astrochemistry of PAHs and PANHs and astronomical searches for such species at radio through (sub)mm wavelengths.